Reports 1-1 of 1 Clear search Modify search
koki.okutomi - 19:12 Tuesday 03 March 2020 (13304) Print this report
ITMY GAS height adjustment: net height not changed that much due to oplev sum's displeasure

[Nakano, Aso, Lucia, Okutomi]

  • We once lifted up the height of the payload by moving the F3 GAS fishing rod.
    • Other fishing rods in F1 and F2 seems not so smooth compared to the F3 GAS in the lifting-up operation.
  • But after the height adjustment raising by 2 mm, we found a strange fluctuation in TM_OPLEV_{LEN, TILT}_SUM.
    • This 2 mm was confirmed in the BF V DC readout.
    • The fluctuation in the oplev sums looks oscillating at a suspension-like frequency such as ~ 1 Hz.
    • At this moment, the amplitude of the sum fluctuation was a few tens of counts in peak to peak.
    • This fluctuation was not resolved even at the height of 1 mm from the initial position.
    • It was difficult to accomplish the interferometer lock in this situation.
  • In order to avoid this strange oplev behavior, we finally recover almost the original height of the payload.
    • To say exactly, the height was slightly raised by ~ 200 um.
    • Now it looks the fluctuation of the oplev sum decreased to a reasonable level like less than 10 counts in its amplitude.
Comments to this report:
shinji.miyoki - 22:48 Tuesday 03 March 2020 (13307) Print this report

Before i and Lucia tried as klog#13208. Is this offset adjustment still alive ?

Search Help